OpenStudy (anonymous):

please help ???:( -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (anonymous):

Throw the '-2x' to the other side in first equation. You get: y = 1 + 2x Plug this equation into second one: -4x + (1 + 2x) = -1 -2x = -2 x = 1 And plug this 'x' value in one of the equations. -4(1) + y = -1 y = 3

OpenStudy (anonymous):

2x + y = 1 4x + 2y = -1 x = 0 x = 2 There is no x value, as there is no solution. x can be any value, as there is an infinite number of solutions.

OpenStudy (anonymous):

Yes. Correct